# DiRW: Path-Aware Digraph Learning for Heterophily

### v1

@@ 4.1.2. A Plug-and-Play Approach
The results offer compelling evidence that the optimized path sampler we have carefully developed for digraphs substantially improves the performance of both RAWGNN and PathNet. The performance gains can be directly linked to the shortcomings of the original PathGNNs, which did not adequately consider the directionality of edges in digraphs, leading to a substantial loss of valuable information. DiRW addresses this shortcoming by balancing the directionality and existence of directed edges into the sampling process, thereby effectively capturing the intricate topological structures present in digraphs. In addition, our walking probabilities which take into account both the topological structure and the node profiles have demonstrated superior performance compared to RAW-GNN and PathNet that solely considered the topology in their walking probability design. By modeling node similarity, DiRW effectively captures one-hop homophily information in heterophilous settings. The successful integration of these two strengths has not only led to improved performance when DiRW is considered as a plug-and-play module with other PathGNNs but also underscores the adaptability and robustness of our approach.
| Models | CiteSeer | WikiCS | Chameleon | Rating |
| w/o Dir | 65.111.77 | 79.650.18 | 42.471.07 | 45.590.16 |
| w/o Topo | 64.260.17 | 79.370.13 | 38.972.64 | 45.790.18 |
| w/o Feat | 65.771.28 | 79.270.13 | 41.651.73 | 45.620.25 |
| w/o Att | 65.020.75 | 79.260.32 | 39.592.23 | 42.550.39 |
| DiRW-Gate | 65.650.84 | 79.520.28 | 38.763.56 | 42.680.15 |
| DiRW-JK | 63.881.02 | 79.310.25 | 40.213.82 | 42.740.33 |
| DiRW | 65.881.41 | 79.870.08 | 43.920.23 | 46.130.39 |
4.2. Ablation Study
To answer Q2, we have conducted a series of ablation studies to ascertain the individual contributions of the six core components that constitute the backbone of DiRW. Starting with the direction-aware path sampler, we set the walk direction coefficient to 1 to simulate a sampling scenario without directionality (w/o Dir). Next, we conducted separate ablations for the topological-based (w/o Topo) and feature-based (w/o Feat) walk probabilities to understand the relative importance of node profiles and topological structure in guiding the walks. Additionally, the ablation experiment for our node-wise attention mechanisms is conducted, which is replaced by directly averaging path embeddings (w/o Att), Gate attention proposed by (Ahmad et al., 2021) (DiRW-Gate) and JK attention proposed by (Xu et al., 2018) (DiRW-JK). The outcomes of these ablation studies are detailed in Tab. 3. Furthermore, to evaluate the adaptive walk length and numbers, we replace them with fixed hyperparameters and depict the outcomes in Fig. 4. We measured the performance and running time under varying walk lengths and number, while keeping other hyperparameters consistent with DiRW to ensure a fair comparison. We draw the following conclusions from the results:
Direction-aware path sampler. By effectively balancing the directionality and existence of edges in digraphs, the direction-aware path sampler has significantly enhanced the model’s capacity to capture the intricate topological structures of digraphs, thereby resulting in an average performance improvement of 1.5%.
Topological walk probability. Designed to facilitate exploration of high-order homophilous neighbors, the topological-based walk probabilities have demonstrated a marked advantage in detecting homophilous patterns within the digraphs, contributing to notable enhancements in the model’s predictive performance.
Feature-based walk probability. By harnessing node profiles to guide the exploration towards nodes with strong homophilous ties, the feature-based walk probability has successfully driven an average performance increase of 1.9%, highlighting the critical role of feature information in guiding walk preference.
Node-wise Learnable Path Aggregator. By employing our nuanced attention mechanism to discern the salience of different paths, DiRW adeptly differentiates between effective and ineffective sampling sequences, thereby distilling high-quality node embeddings and results in an average performance boost of 5.36%. The node-wise attention mechanism excels by integrating a dual MLP, which is adept at capturing complex and subtle nuances in the path embeddings as compared to the single MLP used in the JK and Gate attention mechanisms. The dual-layer MLP not only processes the initial linear transformation of the path embeddings but also applies a subsequent layer that refines the feature representation.
Homophily entropy-based personalized walk length. The personalized walk length has proven to be a boon for DiRW performance, outclassing any model reliant on a static walk length. Traditional approaches with longer fixed walk length often necessitate substantial pre-processing time and yield sub-optimal performances. In contrast, DiRW achieves a commendable efficiency and optimal effectiveness. By leveraging homophily entropy to assess the path quality, DiRW is adept at identifying and sampling high-quality sequences promptly, thus streamlining the sampling process and ensuring that the model operates at an optimal level.
Heuristic adaptive walk number. Echoing the success of the personalized walk length, the heuristic adaptive walk number also demonstrates a significant performance improvement over models that utilize a rigid number of walks. This superior efficiency and effectiveness stems from our systematic evaluation of the information richness within sampled walks, allowing the model to sample a informative set of walk sequences in a shorter span.
4.3. Efficiency Analysis
To address Q3, we conducted experiments to evaluate the efficiency of DiRW. We selected an spatial-based baseline ADPA and two spectral-based baselines Magnet and LightDiC to benchmark their performance and efficiency. The metric for efficiency was established based on the total time required to run each model for five times, with the results graphically depicted in Fig. 5.
From the visual representation, it is evident that the superiority of DiRW lies in both predictive accuracy and operational efficiency. Despite the variability in the optimal adaptive walk length and walk numbers across different datasets, which might introduce some instability in computational time, DiRW maintains an acceptable time cost and consistently outperforms ADPA. The result reveals that ADPA is consistently the most time-consuming model. This is largely due to its intricate architecture, featuring hierarchical attention mechanisms and multiple convolution layers, which increases its computational complexity and resource requirements.
Significantly, the majority of DiRW’s running time is consumed by the sampling stage, which can be a bottleneck in the overall training process. However, by pre-processing this sampling stage and storing the sequences, which is similar to the approach employed by PathNet, we can considerably enhance the efficiency of the training pipeline. Furthermore, DiRW’s rapid learning mechanism is designed to leverage these pre-processed sequences effectively, facilitating a more streamlined and accelerated training process.
4.4. Path Quality Analysis
To answer Q4, we conducted an in-depth analysis comparing the path quality sampled by our optimized path sampler with those generated by the DFS and BFS utilized in RAWGNN. This analysis is examined by the homophily entropy according to Eq. (8).
The experimental outcomes in Fig. 6 demonstrate that the paths sampled by DiRW consistently exhibit lower homophily entropy compared to those produced by DFS and BFS. This superiority is primarily attributed to the integrated consideration of both topological structure and node profiles in walk probability mechanism, coupled with the adaptive walk length guided by homophily entropy. Furthermore, the DFS approach demonstrated the highest homophily entropy in most scenarios. This can be attributed to the fact that BFS, with its mechanism of exploring the immediate neighborhood, aligns well with the homophily assumption. However, in scenarios characterized by strong heterophily like actor, DFS outperforms BFS in path sampling quality. Our DiRW model achieves a harmonious balance between these two approaches, leveraging the strengths of both while mitigating their respective weaknesses.
5. Conclusion and Future Work
In this work, we have underscored the necessity of DiGNNs in the context of modern graph-structured data and proposed DiPathGNN mechanism to capture the inherent homophilous information in digraphs. Through empirical analysis, we have exposed the limitations of current PathGNNs, particularly in their lack of edge direction modeling and coarse-grained sampling strategy. To address these challenges, we propose DiRW, a novel path-aware digraph learning approach. DiRW is distinguished by its direction-aware path sampler, which is meticulously fine-tuned based on walk probability, length, and numbers in a weight-free approach. Furthermore, DiRW employs a node-wise learnable path aggregator to formulate nuanced node representations. Experiments demonstrate that DiRW achieves SOTA performance in node- and link-level tasks, especially for topologically heterophilous scenarios.
Looking ahead, there are several promising directions for future research. First, we aim to explore more efficient sampling strategies that can further enhance the training effectiveness. Furthermore, developing adaptive sampling methods that dynamically adjust based on the characteristics of the graph could potentially improve the performance and scalability of our approach. Lastly, designing better learning mechanisms to capture the complex relationships between nodes in digraphs could lead to more accurate predictions. These future works will further enhance our understanding of digraph learning and potentially lead to breakthroughs in how we model and analyze complex graph-structured data.
Appendix A Outline
- A.1:
Model implementations in Empirical Study.
- A.2:
Datasets Description.
- A.3:
Baselines Description.
- A.4:
Hyperparameter Settings.
- A.5:
Experimental Environment.
- A.6:
DiRW Performance in Link Prediction.
- A.7:
Sensitivity Analysis.
- A.8:
Comparison of DiRW with Other PathGNNs.
- A.9:
PathGNN and Graph Attention.
A.1. Model implementations in Empirical Study
Inspired by the concepts of node homophily (Pei et al., 2020) and edge homophily (Zhu et al., 2020), which are pivotal in understanding the homophily of graphs, we introduce a refined definition of the node homophily ratio for each node:
where represents the set of neighbors of node , and denotes the label of node . The node homophily ratio serves as a metric for the homophily of the local neighborhood around node . A higher value of indicates stronger homophily, meaning node is more similar to its neighbors in terms of the given label. In experiment, we select the top 50% most homophilous nodes as homophily nodes, while the remaining nodes as heterophily nodes. When investigating the impact of different walk lengths on the model, we fixed the number of walks at . Similarly, we fixed the walk length at when exploring the effect of walk numbers.
We utilize the undirected SRW as our sampling strategy. Specifically, each node randomly selects a neighbor from its in-neighbors and out-neighbors as the next target. For each node in the graph, we specify the same number of walks and the same walk length . Once each walk sequence is collected, we concatenate the features of every node in the sequence. Suppose the walk sequence is , where represents the -th node in the walk sequence. For each walk sequence, we concatenate the node features to obtain the path feature vector:
where denotes the feature vector of node . We then pass this path feature vector through a MLP to obtain the path embedding:
For each node, we average the embeddings of its different paths to obtain the node embedding. Let the set of path embeddings for node be . The embedding for node is:
Finally, we pass the node embedding through another MLP for the node classification task:
By leveraging this approach, we effectively utilize path information to enhance node representations, achieving better performance in node classification tasks.
A.2. Datasets Description
We evaluate the performance of DiRW under 9 benchmark datasets, covering both homophilous and heterophilous graphs. We have documented the detailed information of these datasets in the Tab. 4. Below, we introduce the descriptions of these benchmark datasets.
| Actor | 7,600 | 26,659 | 932 | 5 | 48%/32%/20% | actor network | |
| Rating | 24,492 | 93,050 | 300 | 5 | 50%/25%/25% | rating network |
CoraML and CiteSeer (Bojchevski and Günnemann, 2018) are two academic citation networks, where individual academic papers are represented as nodes, connected by edges that signify citations made between these papers. Each node is characterized by a binary word vector that encapsulates its textual content, while the class labels correspond to the subject areas to which the papers pertain.
WikiCS (Mernyei and Cangea, 2020) is a dataset derived from Wikipedia. It comprises nodes that represent articles within the computer science domain, with edges established based on the hyperlinks that connect these articles. The dataset is organized into 10 distinct classes, each representing a different subfield of computer science. The features are generated from the text content of the respective articles by averaging pre-trained GloVe word embeddings (Pennington et al., 2014), yielding a 300-dimensional vector for each node.
Amazon-Computers (Shchur et al., 2018) is a co-purchase network from Amazon that encapsulates consumer behavior through a graph structure. In this graph, nodes correspond to computer products, and edges denote the instances where products are frequently bought together by customers. The features for each product are derived from the text of product reviews, represented as bag-of-words vectors that capture the essence of customer feedback.
Ogbn-arxiv (Hu et al., 2020) is a citation network dataset indexed by the MAG (Wang et al., 2020), which is characterized by the word embeddings found within the titles and abstracts of academic papers. The word embeddings are generated through the application of the skip-gram model, which is for learning vector representations of words.
Ogbn-products (Hu et al., 2020) is a co-purchasing network where nodes represent products and edges represent two products are frequently bought together. Node features are generated by extracting bag-of-words features from the product descriptions.
Chameleon (Rozemberczki et al., 2021) is a network derived from English Wikipedia, where nodes represent articles and edges indicate mutual backlinks between them. Each node is characterized by a feature set that reflects the presence of specific nouns.
Actor (Pei et al., 2020) is an actor co-occurrence network where each node symbolizes an actor and the edges indicate instances where actors are mentioned together on Wikipedia pages. The node features in this network are constructed as bag-of-words vectors, which are created by extracting keywords from the Wikipedia pages dedicated to each actor. These features are then sorted into five different categories, each corresponding to the specific terms present on an actor’s Wikipedia page.
Rating (Platonov et al., 2023) is a dataset that originates from the Amazon product co-purchasing network metadata found within the SNAP (Leskovec and Krevl, 2014). In this network, nodes correspond to different products, and edges are established between items that are commonly purchased in conjunction with one another. The primary task is to forecast the average rating awarded by reviewers, which is divided into five distinct classes. The node features are formulated from the mean FastText embeddings (Grave et al., 2018) of the words present in the product descriptions, providing a textual representation of each product.
A.3. Baselines Description
Our experiment leverages diverse GNNs as comparative benchmarks, which can be categorized as follows: (i) traditional undirected approaches: GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018); (ii) directed spatial methods: DGCN (Tong et al., 2020b), NSTE (Kollias et al., 2022), DIMPA (He et al., 2022), Dir-GNN (Rossi et al., 2023), ADPA (Sun et al., 2023); (iii) directed spectral methods: DiGCN (Tong et al., 2020a) and its variant DiGCNappr, MagNet (Zhang et al., 2021a), MGC (Zhang et al., 2021b), LightDiC (Li et al., 2024a); (iiii) PathGNNs: RAW-GNN (Jin et al., 2022) and PathNet (Sun et al., 2022). To minimize randomness and ensure fair comparisons, we repeated each experiment 10 times to obtain unbiased performances. Moreover, we transform the digraph to undirected graph and feed it into undirected GNNs. The main features of each baseline model are as follows:
GCN (Kipf and Welling, 2017): GCN operates on the principle of a localized first-order spectral graph convolution, which serves as an approximation. It is designed to scale efficiently with the number of edges in the graph. The model processes information through hidden layers that synthesize the local neighborhood structure and individual node features, thereby learning representations that are pertinent to the graph’s topology and node attributes.
GAT (Veličković et al., 2018): GAT employs attention mechanisms to dynamically assess the significance of neighboring nodes during the aggregation of information. This approach allows for the flexible assignment of varying influence to different nodes within a neighborhood, without the need for predefined knowledge of the graph’s structure.
DGCN (Tong et al., 2020b): DGCN pioneers a novel message-passing scheme that takes into account both the first and second nearest neighbors’ proximity. This approach results in the development of aggregators that are tailored to differentiate between incoming and outgoing edges, utilizing two distinct sets of parameters that are learned through the training process.
DiGCN (Tong et al., 2020a): Recognizing the intrinsic relationship between the graph Laplacian and the steady-state distributions of PageRank, DiGCN introduces a theoretical advancement that adapts personalized PageRank to formulate a genuinely symmetric representation of the directed graph Laplacian. Additionally, DiGCN leverages the closeness of both first and second-order neighbors to augment the model’s receptive field.
DiGCNappr (Tong et al., 2020a): This variant of DiGCN incorporates an efficient approximation for personalized PageRank, streamlined by the absence of inception blocks. It represents a generalized form of GCN that is grounded in the directed graph Laplacian, offering a more nuanced approach to graph convolution for directed networks.
DIMPA (He et al., 2022): In the DIMPA model, the source and target nodes are distinctly represented. It gathers neighborhood information across multiple hops within each layer to expand the model’s receptive field. Furthermore, DIMPA employs a weighted averaging mechanism for the aggregated multi-hop neighborhood data, which allows it to effectively encapsulate the local network structure.
NSTE (Kollias et al., 2022): NSTE draws inspiration from the 1-WL graph isomorphism test, utilizing two distinct sets of weights to encode the source and target nodes independently. Subsequently, the model refines the weights used for information aggregation through a parameterized feature propagation process.
Dir-GNN (Rossi et al., 2023): The Dir-GNN framework is specifically crafted to excel in heterophilous scenarios. It effectively manages the directionality of edges through distinct aggregation processes for incoming and outgoing edges. This approach has been shown to rival the expressiveness of the directed Weisfeiler-Lehman test.
ADPA (Sun et al., 2023): ADPA is distinguished by its adaptive exploration of appropriate neighborhood operators of directed -order for propagation without the need for edge weights. It also incorporates a two-tiered system of node-adaptive attention mechanisms, which are designed to refine and optimize the representation of nodes within the graph. This adaptive methodology allows ADPA to dynamically select the most suitable operators for the given graph structure, leading to enhanced node representation capabilities.
MagNet (Zhang et al., 2021a): MagNet employs the mathematical framework of complex numbers to capture the nuances of directed information flow. It introduces a novel spectral approach to GNNs for digraphs, leveraging the concept of a complex Hermitian matrix termed the magnetic Laplacian. To further enhance predictive accuracy, MagNet incorporates additional parameters that independently merge the real and imaginary components of the filter signals.
MGC (Zhang et al., 2021b): MGC introduces the magnetic Laplacian, a discrete operator that incorporates a magnetic field effect. It also employs a truncated form of the PageRank algorithm, referred to as Linear-Rank, to construct filters tailored for both homogeneous and heterogeneous graphs, with a low-pass and high-pass filter, respectively.
LightDiC (Li et al., 2024a): LightDiC is an innovative and scalable adaptation of digraph convolution that is founded on the magnetic Laplacian. It performs all topological computations during an offline pre-processing phase, which showcases its scalability. It allows for the downstream tasks to be trained in an independent manner, without the computational burden with recursive processes.
RAW-GNN (Jin et al., 2022): RAW-GNN leverages the Node2Vec algorithm to emulate both breadth-first search (BFS) and depth-first search (DFS) traversal techniques. This dual-simulation approach enables the model to encapsulate a comprehensive spectrum of information within the graph, including both the tendencies of nodes to connect with similar and dissimilar neighbors.
PathNet (Sun et al., 2022): PathNet is distinguished by its use of a maximal entropy-based random walk strategy. By optimizing for maximal entropy, PathNet ensures a diverse and comprehensive exploration of the graph’s topology, leading to more robust and informative node representations.
A.4. Hyperparameter Settings
A.5. Experimental Environment
Our experiments are conducted on the machine with Intel(R) Xeon(R) Platinum 8468V, and NVIDIA H800 PCIe and CUDA 12.2. The operating system is Ubuntu 20.04.6. As for the software, we use Python 3.8 and Pytorch 2.2.1 for implementation.
A.6. DiRW Performance in Link Prediction
@@
In addition to the node classification task, we have conducted a thorough evaluation of DiRW’s performance in the link direction prediction task, with the results presented in Tab. 5. We directly leverage the node embeddings obtained from Eq. (LABEL:eq:_Attention) and concatenate them to form edge embeddings, with Eq. (LABEL:eq:_classifier) obviated. This concatenated vector serves as a comprehensive representation of the edges, capturing the characteristics between connected nodes.
@@
The experimental outcomes provide compelling evidence that DiRW excels not only in categorizing nodes but also in predicting the direction of edges, outperforming the leading Dir-GNN by an average of . The efficacy of DiRW in link prediction is a testament to its ability to capture the subtleties of graph structure and the nuanced relationships between nodes. This capability stems from the model’s sophisticated path sampling strategies and its attention-based aggregation mechanism, which together allow for a deep understanding of the digraph’s topology and node features.
@@
It is important to note that traditional undirected GNNs have exhibited subpar performance in this task. The primary reason for this deficit lies in the task’s requirement for precise capture of the directional information of edges between nodes, a nuance that conventional undirected methodologies fail to address adequately. In contrast, DiRW’s design explicitly accounts for edge directionality, providing it with a distinct advantage in tasks where the direction of relationships is pivotal. This comparative analysis underscores the importance of adopting models that can accommodate the directed nature of graphs, thereby offering a more accurate representation of the underlying structure and relationships within the data.
A.7. Sensitivity Analysis

### v3

@@ 4.2.2. A Plug-and-Play Approach
The results provide strong evidence that our optimized path sampler significantly enhances the performance of both RAWGNN and PathNet. The performance gains stem from the shortcomings of the original PathGNNs, which fail to consider adequately the directionality of edges in digraphs, leading to a substantial loss of valuable information. DiRW corrects it by balancing the edge directionality and existence into its sampling process, effectively capturing the intricate topological structures inherent in digraphs. Moreover, our walking probabilities, which account for topological structure and node profiles, outperform those of RAWGNN and PathNet, which focus solely on topology. By modeling node similarity, DiRW effectively captures one-hop homophily in heterophilous contexts. This integration not only enhances performance when DiRW is used as a plug-and-play module with other PathGNNs but also highlights the adaptability and robustness of our approach.
4.2.3. Performance in Link Prediction
@@
We extend the evaluation of DiRW through a multi-task framework, investigating both node classification and directional link prediction capabilities, with the results presented in Tab. 4. We directly leverage the node embeddings derived from Eq. (LABEL:eq:_Attention) and concatenate them to form discriminative edge embeddings, thereby bypassing the need for additional node classifier as specified in Eq. (11). This concatenated vector serves as a comprehensive representation of the edges, preserving the geometric relationships encoded during the walk aggregation phase while maintaining architectural simplicity.
@@
The experimental outcomes provide compelling evidence that DiRW achieves an average performance gain of over the second leading model Dir-GNN. The efficacy of DiRW in link prediction is a testament to its ability to capture the subtleties of graph structure and the nuanced relationships between nodes. This capability stems from the model’s sophisticated path sampling strategies and attention-based aggregation mechanism, which allow for a deep understanding of the digraph’s topology and node features.
@@
It is important to note that traditional undirected GNNs have exhibited subpar performance and suffer from structural isomorphism issues in direction-sensitive tasks. The primary reason for this deficit lies in the direction-agnostic message-passing in undirected architectures, which fails to discriminate between () and () edge semantics. In contrast, DiRW’s design explicitly accounts for edge directionality, providing it with a distinct advantage in tasks where the direction of relationships is pivotal. This comparative analysis underscores the importance of adopting models that can accommodate the directed nature of graphs, thereby offering a more accurate representation of the underlying structure and relationships within the data.
4.3. Ablation Study
| Models | CiteSeer | WikiCS | Chameleon | Rating |
| w/o Dir | 65.111.77 | 79.650.18 | 42.471.07 | 45.590.16 |
| w/o Topo | 64.260.17 | 79.370.13 | 38.972.64 | 45.790.18 |
| w/o Feat | 65.771.28 | 79.270.13 | 41.651.73 | 45.620.25 |
| w/o Att | 65.020.75 | 79.260.32 | 39.592.23 | 42.550.39 |
| DiRW-Gate | 65.650.84 | 79.520.28 | 38.763.56 | 42.680.15 |
| DiRW-JK | 63.881.02 | 79.310.25 | 40.213.82 | 42.740.33 |
| DiRW | 65.881.41 | 79.870.08 | 43.920.23 | 46.130.39 |
To answer Q2, we have conducted a series of ablation studies to ascertain the individual contributions of the five core components that constitute the backbone of DiRW.
We began with the direction-aware path sampler by setting the walk direction coefficient to , simulating a scenario without directionality (w/o Dir). Then, we conducted separate ablations for the topology-based (w/o Topo) and feature-based (w/o Feat) walk probabilities to assess the relative importance of node profiles and topological structure in guiding the walk. Additionally, we conducted an ablation experiment on our node-wise attention mechanisms, replacing them with direct averaging of path embeddings (w/o Att), Gate attention (Ahmad et al., 2021) (DiRW-Gate), and JK attention (Xu et al., 2018) (DiRW-JK). The outcomes are detailed in Tab. 5.
Furthermore, to assess the adaptive walk length and numbers, we replace them with predefined and fixed hyperparameters and depict the outcomes in Fig. 4. We measured the performance and running time across varying walk lengths and numbers while keeping other hyperparameters consistent with DiRW. From these results, we draw the following conclusions:
Direction-aware path sampler. By effectively balancing the direction and existence of edges, the direction-aware path sampler has significantly enhanced the model’s capacity to capture the intricate topological structures of digraphs, leading to an average performance improvement of 1.5%.
Multi-order walk probability. The topology-based high-order walk probability shows a clear advantage in detecting homophilous patterns, leading to significant improvements in predictive performance. Moreover, the feature-based one-order walk probability achieves an average performance increase of 1.9%, underscoring feature information’s importance in guiding walk preferences.
Node-wise learnable path aggregator. By utilizing our attention mechanism, DiRW effectively distinguishes between productive and unproductive sampling sequences, leading to high-quality node embeddings and an average performance boost of 5.36%. The node-wise attention mechanism employs a dual MLP, which captures complex nuances in path embeddings more effectively than the single MLP used in the JK and Gate attention mechanisms.
Homophily entropy-based personalized walk length. The personalized walk length greatly enhances DiRW’s performance, outperforming static walk length models that require significant pre-processing and often yield sub-optimal results. In contrast, DiRW leverages the homophily entropy to sample high-quality paths, streamlining the sampling process for optimal effectiveness.
Adaptive walk number. Echoing the success of the personalized walk length, the adaptive walk number also demonstrates a significant performance improvement over models that utilize a rigid walk number. It comes from our systematic evaluation of the information richness in sampling more informative walk sequences.
4.4. Efficiency Analysis
To address Q3, we conducted experiments to evaluate the efficiency of DiRW. We selected a spatial-based baseline, ADPA, and two spectral-based baselines, Magnet and LightDiC, to benchmark their performance and efficiency. The metric for efficiency was established based on the total time required to run each model five times, with the results graphically depicted in Fig. 5.
From the visual representation, it is evident that the superiority of DiRW lies in both predictive accuracy and operational efficiency. Although optimal adaptive walk lengths and numbers vary between datasets, causing some fluctuations in computational time, DiRW achieves the best performance and efficiency due to its lightweight learning mechanism. ADPA is consistently the most time-consuming model, largely due to its intricate architecture, featuring hierarchical attention mechanisms and multiple convolution layers, which increases its computational and resource complexity.
Notably, the sampling stage constitutes the primary computational bottleneck in DiRW’s workflow, accounting for the majority of its running time. By adopting a pre-processing strategy analogous to PathNet, where we precompute and store sampled sequences, we can substantially optimize the training pipeline’s efficiency. This preparatory approach synergizes effectively with DiRW’s dedicated rapid learning mechanism, facilitating more accelerated training.
4.5. Sensitivity Analysis